OpenStudy (anonymous):

Determine which order pair is a solution of y= 2x - 3? (-2,7) (0,3) (-4,-11) (5,-7)

OpenStudy (dan815):

the points which satisfy that equation

OpenStudy (butterflydreamer):

You want to just sub in your points into the equation "y = 2x - 3" and check if they're correct. So for example, substitute in the first point (-2,7) into the equation "y = 2x - 3" This means your x = -2 and y = 7 Therefore when you sub it into the equation you get y = 2x - 3 7 = 2 (-2) - 3 7 = -4 - 3 7 = -7 Does 7 equal to -7? No! So then this is incorrect and therefore (-2,7) is NOT a solution :) Basically sub in the point and you want the LHS = RHS

OpenStudy (anonymous):

ok so whats the answer

OpenStudy (butterflydreamer):

lmaonaise work it out. Don't be lazy LOL